OpenStudy (anonymous):

I need help solving 6x^2+13x-28 by factoring??

OpenStudy (anonymous):

solve this equation: 6x^2+13x-28 = 0 ???

OpenStudy (anonymous):

if yes, then: 6x^2+13x-28 = 0 (2x+7)(3x-4) = 0 then set each factor = 0 and solve.